janet found 30 coins in her car. all of the coins were eighter nickels or dimes. if the total value was $2.10, find the number of each type of coin?
Answer by checkley77(12844)

D+N=30 OR D=30-N
.10D+.05N=2.10
.10(30-N)+.05N=2.10
3-.10N+.05N=2.10
-.05N=2.10-3
.05N=-.90
N=.9/.05
N=18 NICKLES.
30-18=12 DIMES.
PROOF:
.10*12+.05*18=2.10
1.20+90=2.10
2.10=2.10